The market demand for wheat is Q = 100 - 2p + 1pb + 2Y. If the price of wheat, p, is $2,

and the price of barley, pb, is $3, and income, Y, is $1000, what is the income elasticity of

wheat.


1
Expert's answer
2021-05-24T13:53:44-0400

"Q = 100 - 2p +1pb + 2Y"


By putting all given values in Q we get


"Q = 100 - (2 \u00d7 2) + (1 \u00d7 3) + (2 \u00d7 1000)"


"Q = 100 - 4 + 3 + 2000"


"Q = 2099"


Income elasticity of wheat (Ey) = "\frac {\Delta Q} {\Delta Y} ×\frac {Y} {Q}"


"Ey = 2 ×(\frac {1000} {2099})"


"Ey = 0.95"


Therefore, here the income elasticity of wheat is 0.95